OpenStudy (anonymous):

If y=e^-x * cos(2x), then dy/dx is what?

OpenStudy (dumbcow):

use product rule \[(fg)' = f'g + fg'\] \[f' = -e^{-x}\] \[g = -2 \sin(2x)\]

OpenStudy (anonymous):

Use the product rule (uv)' = u'v + v'u \[u= e^{-x} \\ v =\cos(2x) \]

OpenStudy (anonymous):

\[ y=e^{-x} \cos(2x)\\ y'=- e^{-x} \cos(2x) -2 e^{-x} \sin(2x) \]